southwest airlines plane: A southwest airlines plane can fly 455mph against the wind and 515mph when it flies with thw wind. find the effect of the wind and the groundspeed of the airplane.
Answer by jorel1380(3719)

455+515/2=485
485-455=30
515-485=30
The plane's speed in still air is 485 mph. The wind speed is 30 mph..